# High-dimensional limit theorems for SGD: Momentum and Adaptive Step-sizes

### v1

@@ 3.2. Single Index Models
Similar to the previous example, we determine the evolution of the pair of summary statistics $\boldsymbol{u}_{n}=(m,r_{\perp}^{2})$. The scaling limit for this problem was analyzed for online SGD by [saad1995dynamics, saad1995line, goldt2019dynamics, veiga2022phase, single_index]. The critical scaling is determined by the step-size $\delta_{n}=c_{\delta}/n$.
@@
(9) $\text{d}m$ $=\frac{-2}{(1-\beta)}\mathbb{E}\left[a_{1}f^{\prime}\left(a_{1}m+a_{2}r_{\perp}\right)\left(f\left(a_{1}m+a_{2}r_{\perp}\right)-f(a_{1})\right)\right]\,\text{d}t,$
$\text{d}r_{\perp}^{2}$ $=\frac{-4}{(1-\beta)}\bigg\{\mathbb{E}\left[a_{2}r_{\perp}f^{\prime}\left(a_{1}m+a_{2}r_{\perp}\right)\left(f\left(a_{1}m+a_{2}r_{\perp}\right)-f(a_{1})\right)\right]$
(10) $\quad\quad-\left.\frac{c_{\delta}}{1-\beta}\mathbb{E}\left[f^{\prime}\left(a_{1}m+a_{2}r_{\perp}\right)^{2}\left((f\left(a_{1}m+a_{2}r_{\perp}\right)-f(a_{1}))^{2}+\sigma^{2}\right)\right]\right\}\,\text{d}t$
@@
(12) $\text{d}r_{\perp}^{2}$ $=-2\left\{r_{\perp}\mathbb{E}[a_{2}\cdot\text{sgn}(f^{\prime}(ma_{1}+ra_{2})(f(ma_{1}+ra_{2})-f(a_{1})+\epsilon))]-\frac{c_{\delta}}{2}\right\}\,\text{d}t$
@@
In what follows, we consider $f(x)=x^{k}$ for $k\geq 1$. This family of link functions includes well-studied examples such as phase retrieval ($k=2$). In the small noise regime $\sigma^{2}\to 0$, the population optimum $(m,r^{2})=(1,0)$ is a fixed point for SGD-M (see Proposition 3.8), but not for SGD-U. Despite this, we show that SGD-U is able to converge to fixed points that are still close to the optimum in settings where SGD-M diverges (i.e. $\text{d}r_{\perp}^{2}>0,\,\forall t$).
@@
$\text{d}m$ $=-\frac{12}{1-\beta}m(R^{2}-1)\,\text{d}t,\quad\text{d}r^{2}=-\frac{4}{1-\beta}\left(2r^{2}(3R-1)-\frac{4c_{\delta}}{1-\beta}\left(P(m,r)+\sigma^{2}R^{2}\right)\right)\,\text{d}t$
@@
    For $c_{\delta}>(1-\beta)/12$, we have a fixed point at $(m,r^{2})=(0,0)$ and additional fixed points $(m,r^{2})=(0,r^{2}_{p})$ for $r^{2}_{p}>0$ where $r_{p}^{2}$ are roots to the quadratic equation
@@
The proof for both results is provided in Section 6.2. These results inform us of the inherent difficulty for the SGD-M (and consequently online SGD) dynamics to converge to fixed points near the population minimum. To attain these fixed points, we require taking a smaller $c_{\delta}$, thus increasing number of iterations to run the algorithm as per our discussion in Section 3.1. We capture this phenomena more generally in the small noise limit next.
@@ 3.2.1. Small Noise $\sigma^{2}\to 0$ Limit
For SGD-M, increasing the degree of the monomial link restricts the admissible values of $c_{\delta}$ for which the dynamics converge to fixed points away from $m=0$. In particular, one can show (e.g. see Section 6.2) that the rate for which the upper bound decreases is of the order $k^{-k-2}$. In particular, we observe that even though $(1,0)$ is a fixed point for any $c_{\delta}>0$ when $\sigma^{2}=0$, it is not stable unless $c_{\delta}$ satisfies (13). We next determine the dynamics for a large collection of link functions for SGD-U in the small-noise regime.
@@
These dynamics yield a unique fixed point $(m,r^{2})=(1,c_{\delta}/(2\sqrt{2/\pi}))$.
This can be viewed as a type of “universality” result in the link function. For odd-powered monomials, (13) illustrates how $c_{\delta}$ must be super-exponentially small in $k$ for the dynamics of SGD-M to converge to a fixed point about $m=1$. However, in Proposition 3.9, we see that regardless of the degree, the dynamics under SGD-U always converge to a fixed point along $m=1$ with $r^{2}>0$. Setting $c_{\delta}$ arbitrarily small thus recovers fixed points near the population optimum. We can view this as a rigorous justification for how using normalized gradients, as in SGD-U, allows for choosing larger step-sizes for which the learning dynamics remain reasonable. This example validates some of the heuristics and intuition behind using unit-norm gradients to control exploding and vanishing gradients, causing instability for SGD [mikolov2012statistical, pascanu2013difficulty].
@@ Acknowledgments
Let $\tau_{K}^{n}$ denote the exit time of the interpolated process $\boldsymbol{u}_{n}(t)$ from the compact set $E_{K}^{n}$. Furthermore, let $E^{*}_{K,n}:=\boldsymbol{u}_{n}^{-1}(E^{n}_{K})$ denote the corresponding pre-image, and $L^{\infty}_{K,n}:=L^{\infty}(E^{*}_{K,n})$ the associated space of localized bounded functions. The subsequent analysis relies on several technical lemmas, which hold under the $\delta_{n}$-closability and -localizability assumptions detailed in Definition 2.1, 2.2. We assume $\boldsymbol{p}_{0}=0$ unless stated otherwise.
@@
By consolidating Lemmas 4.5, 4.6, and 4.7, we construct the auxiliary process $\boldsymbol{z}_{n}$ as follows. Consider the $j$-th component $z^{j}$, the discrete-time process is defined by the Doob decomposition:
@@
The remainder of the proof closely follows [benarous2024high, Theorem 2.3]; we provide the details for completeness. Let $\boldsymbol{z}_{n}(s)$ be the linear interpolation of the discrete process $(\boldsymbol{z}_{\lfloor s/\delta\rfloor})$ given in (14). We decompose this process as
@@
The summation ranges over $\ell$ from $\lfloor s/\delta\rfloor+1$ to $\lfloor t/\delta\rfloor$, stopped at $\tau_{K}^{n}/\delta$. The bound follows from the continuity of $\boldsymbol{h}$ as stipulated by $\delta$-closability, i.e. Definition 2.2.
@@
By $\delta$-closability, we ensure that as $n\to\infty$,
@@
By Itô’s formula for continuous semimartingales [stroock2006multidimensional, Theorem 6.3.4], for any $f\in C_{0}^{\infty}(\mathbb{R}^{k})$, the process $f(\boldsymbol{z}^{K}_{t})-\int_{0}^{t}\mathsf{L}f(\boldsymbol{z}^{K}_{s})ds$ is a martingale, where $\mathsf{L}$ is the infinitesimal generator
@@
stopped at the exit time $\tau_{K}$. By employing a standard localization argument (see, e.g. [stroock2006multidimensional, Chapter 11]), we conclude that every limit point $\boldsymbol{z}(t)$ of $\boldsymbol{z}_{n}(t)$ is a solution to the limiting SDE. ∎
4.1. Proof of Lemmas 4.1-4.3

### v2

@@ 3.2. Single Index Models
Similar to the previous example, we determine the evolution of the pair of summary statistics $\boldsymbol{u}_{n}=(m,r^{2})$. The scaling limit for this problem was analyzed for online SGD by [47, 46, 17, 54, 43]. The critical scaling is determined by the step-size $\delta_{n}=c_{\delta}/n$.
@@
(9) $\text{d}m$ $=\frac{-2}{(1-\beta)}\mathbb{E}\left[a_{1}f^{\prime}\left(a_{1}m+a_{2}r\right)\left(f\left(a_{1}m+a_{2}r\right)-f(a_{1})\right)\right]\,\text{d}t,$
$\text{d}r^{2}$ $=\frac{-4}{(1-\beta)}\bigg\{\mathbb{E}\left[a_{2}rf^{\prime}\left(a_{1}m+a_{2}r\right)\left(f\left(a_{1}m+a_{2}r\right)-f(a_{1})\right)\right]$
(10) $\quad\quad-\left.\frac{c_{\delta}}{1-\beta}\mathbb{E}\left[f^{\prime}\left(a_{1}m+a_{2}r\right)^{2}\left((f\left(a_{1}m+a_{2}r\right)-f(a_{1}))^{2}+\sigma^{2}\right)\right]\right\}\,\text{d}t$
@@
(12) $\text{d}r^{2}$ $=-2\left\{r\mathbb{E}[a_{2}\cdot\text{sgn}(f^{\prime}(ma_{1}+ra_{2})(f(ma_{1}+ra_{2})-f(a_{1})+\epsilon))]-\frac{c_{\delta}}{2}\right\}\,\text{d}t$
@@
We verify localizability and closability for both SGD-M and SGD-U as part of the proof of Proposition 3.5 in Section 8. In what follows, we consider $f(x)=x^{k}$ for $k\geq 1$. This family of link functions includes well-studied examples such as phase retrieval ($k=2$). In the small noise regime $\sigma^{2}\to 0$, the population optimum $(m,r^{2})=(1,0)$ is a fixed point for SGD-M (see Proposition 3.8), but not for SGD-U. Despite this, we show that SGD-U is able to converge to fixed points that are still close to the optimum in settings where SGD-M diverges (i.e. $\text{d}r^{2}>0,\,\forall t$).
@@
$\text{d}m$ $=-\frac{12}{1-\beta}m(R^{2}-1)\,\text{d}t,\quad\text{d}r^{2}=-\frac{4}{1-\beta}\left(2r^{2}(3R^{2}-1)-\frac{4c_{\delta}}{1-\beta}\left(P(m,r)+\sigma^{2}R^{2}\right)\right)\,\text{d}t$
@@
    For $c_{\delta}>(1-\beta)/12$, we have a fixed point at $(m,r^{2})=(0,0)$ and additional fixed points $(m,r^{2})=(0,r^{2}_{p})$ for $r^{2}_{p}>0$ where $r_{p}^{2}$ are roots (assuming they exist) to the equation,
@@
The proof for both results is provided in Appendix 8.2. These results inform us of the inherent difficulty for the SGD-M (and consequently online SGD) dynamics to converge to fixed points near the population minimum. To attain these fixed points, we require taking a smaller $c_{\delta}$, thus increasing number of iterations to run the algorithm as per our discussion in Section 3.1. We capture this phenomena more generally in the small noise limit next.
@@ 3.2.1. Small Noise $\sigma^{2}\to 0$ Limit
For SGD-M, increasing the degree of the monomial link restricts the admissible values of $c_{\delta}$ for which the dynamics converge to fixed points away from $m=0$. In particular, one can show (e.g. see Appendix 8.2) that the rate for which the upper bound decreases is of the order $k^{-k-2}$. In particular, we observe that even though $(1,0)$ is a fixed point for any $c_{\delta}>0$ when $\sigma^{2}=0$, it is not stable unless $c_{\delta}$ satisfies (13). We next determine the dynamics for a large collection of link functions for SGD-U in the small-noise regime.
@@
These dynamics yield a unique fixed point at $(m,r^{2})=(1,c_{\delta}^{2}\pi/8)$.
This can be viewed as a type of “universality” result in the link function. For odd-powered monomials, (13) illustrated how $c_{\delta}$ must be super-exponentially small in $k$ for the dynamics of SGD-M to converge to a fixed point about $m=1$. However, in Proposition 3.9, we see that regardless of the degree, the dynamics under SGD-U always converge to a fixed point along $m=1$ with $r^{2}>0$. Setting $c_{\delta}$ arbitrarily small thus recovers fixed points near the population optimum. We can view this as a rigorous justification for how using normalized gradients, as in SGD-U, allows for choosing larger step-sizes for which the learning dynamics remain reasonable. This example validates some of the heuristics and intuition behind using unit-norm gradients to control exploding and vanishing gradients, causing instability for SGD [34, 40].
@@
3.2.2. Diffusive Limits at Fixed Points
Using our results, one can also obtain diffusive limits around appropriate fixed points for rescaled observables, similar to Section 3.1.1. We illustrate this by considering the diffusive limit for phase retrieval at $m=0$ (i.e. at initialization).
Proposition 3.12.
For the link function $f(x)=x^{2}$, The dynamics for the rescaled statistics $\tilde{\boldsymbol{u}}_{n}(t)=(\tilde{m},r^{2})=(\sqrt{n}m,r^{2})$ converges to the solution of,
$\text{d}\tilde{m}$ $=-\frac{12(r^{2}-1)}{1-\beta}\tilde{m}\,\text{d}t+\frac{4r}{1-\beta}\sqrt{c_{\delta}(15r^{4}-18r^{2}+15+\sigma^{2})}\,\text{d}B_{t}$
(14) $\text{d}r^{2}$ $=-\frac{8r^{2}}{1-\beta}\left[3r^{2}-1+\frac{2c_{\delta}(15r^{4}-6r^{2}+3+\sigma^{2})}{1-\beta}\right]\,\text{d}t$
The proof is provided in Appendix 8.3. We note that the evolution of $r^{2}$ is independent of $\tilde{m}$. Furthermore, the equator $\tilde{m}=0$ is unstable when $r^{2}<1$, providing similar insight to the ballistic phase where we observed the emergence of stable fixed points away from the equator for sufficiently small $c_{\delta}$ (which indirectly forces $r^{2}<1$ at such point(s)). One could also obtain a similar diffusive limit for the SGD-U dynamics obtained in Proposition 3.9, however as the fixed point is unique and stable we find that these dynamics may not be as informative.
4. Concluding Remarks
We develop a general theory for the high-dimensional scaling limits of online SGD with Polyak momentum and scalar Markovian pre-conditioning under the effective dynamics framework of [5]. We illustrate our results on two canonical models in high-dimensional inference: Spiked Tensor PCA and Single Index models. For both problems, we consider both the ballistic and diffusive limits of SGD-M and SGD-U, with the latter setting the preconditioner as the inverse gradient norm. In doing so, we provide rigourous examples of how SGD-U is able to stabilize training dynamics and high-dimensional effects leading to more desirable training outcomes as observed empirically. In particular, we illustrate how the effective dynamics framework can sharply capture empirical phenomenon such as exploding/vanishing gradients in the limiting dynamics.
Given these results, there are several interesting directions to explore for future work. Of particular interest is the problem of developing limits of more general scalar preconditioners and develop a systematic analysis comparing and contrasting preconditoning and momentum based methods on broader classes of statistical problems. It would also be interesting to determine the limiting dynamics of popular matrix-based preconditioners [12, 19, 21] and whether they mitigate the high-dimensional corrections.
@@ Acknowledgments
5. Generalization of Theorem 2.3 for Pre-conditioning
Here, we provide a generalized theorem (with adequate conditions) that allows one to prove similar scaling limits in the setting where we adopt scalar Markovian preconditioners for SGD.
Consider the following online learning setup. Suppose we are given an i.i.d. data sequence $\boldsymbol{y}_{1},\boldsymbol{y}_{2},\ldots$ taking values in $\mathcal{Y}\subseteq\mathbb{R}^{d_{n}}$ with law $P_{n}\in\mathcal{M}_{1}(\mathbb{R}^{d_{n}})$ and loss $L:\mathcal{X}_{n}\times\mathcal{Y}_{n}\to\mathbb{R}$ where $\mathcal{X}_{n}\subseteq\mathbb{R}^{p_{n}}$ is the parameter space. Consider online SGD with learning rate $\delta_{n}>0$, momentum rate $\beta\in[0,1)$ and the scalar and Markov preconditioner $\eta_{n}(x,y):\mathcal{X}_{n}\times\mathcal{Y}_{n}\to\mathbb{R}^{+}$, given by:
$\boldsymbol{p}_{\ell}$ $=\beta\boldsymbol{p}_{\ell-1}-\delta_{n}\,\eta_{n}(\boldsymbol{x}_{\ell-1},\boldsymbol{y}_{\ell})\,\nabla L_{n}(\boldsymbol{x}_{\ell-1},\boldsymbol{y}_{\ell})$
$\boldsymbol{x}_{\ell}$ $=\boldsymbol{x}_{\ell-1}+\boldsymbol{p}_{\ell}$
We define the functions
(15) $\nabla\tilde{H}(x,Y)=\eta_{n}(x,Y)\nabla L_{n}(x,Y)-\nabla\tilde{\Phi}(x)\quad\text{ where }\quad\nabla\tilde{\Phi}(x)=\mathbb{E}[\eta_{n}(x,Y)\nabla L_{n}(x,Y)]$
Let $\tilde{V}(x)=\mathbb{E}[\nabla\tilde{H}(x)\otimes\nabla\tilde{H}(x)]$ be the covariance matrix for $\nabla\tilde{H}$ at $x$.
We state two assumptions: $(\delta_{n},\eta_{n})$-localizability and asymptotic-closability.
Definition 5.1.
A tuple $(\boldsymbol{u}_{n},\nabla L_{n},P_{n},\eta_{n},\beta)$ is ($\delta_{n}$, $\eta_{n}$)-localizable with localizing sequence $(E_{K})_{K}$ if there is an exhaustion by compacts $(E_{K})_{K}$ of $\mathbb{R}^{k}$, and constants $C_{K}$ (independent of $n$) such that
  1. (1)
    $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\nabla^{2}u_{i}^{n}\|_{\mathrm{op}}\leq C_{K}\cdot\delta_{n}^{-1/2}$, and $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\nabla^{3}u_{i}^{n}\|_{\mathrm{op}}\leq C_{K}$;
  2. (2)
    $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\nabla\tilde{\Phi}\|\leq C_{K}$, and $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}[\|\nabla\tilde{H}\|^{8}]\leq C_{K}\delta_{n}^{-4}$;
  3. (3)
    $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}[\langle\nabla\tilde{H},\nabla u_{i}^{n}\rangle^{4}]\leq C_{K}\delta_{n}^{-2}$, and
    $\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}[\langle\nabla^{2}u_{i}^{n}(x),\nabla\tilde{H}(x,y)\otimes\nabla\tilde{H}(x,y)-V(x)\rangle^{2}]=o(\delta_{n}^{-3})$, and
    $\sup_{(x_{i})_{1}^{2}\in\boldsymbol{u}_{n}^{-1}(E_{K})}\mathbb{E}_{y_{1}\perp y_{2}}[\langle\nabla^{2}u_{i}^{n}(x_{1}),\nabla\tilde{H}(x_{1},y_{1})\otimes\nabla\tilde{H}(x_{2},y_{2})\rangle^{2}]=o(\delta_{n}^{-3})$ for $\beta>0$
$\tilde{\mathcal{A}}_{n}=\langle\nabla\tilde{\Phi},\nabla\rangle\quad\text{ and }\quad\tilde{\mathcal{L}}_{n}=\frac{1}{2}\langle\tilde{V},\nabla^{2}\rangle$
Definition 5.2.
A family of summary statistics $(\boldsymbol{u}_{n})$ are asymptotically closable for learning and momentum rates $((\delta_{n})_{n},\beta)$, and preconditioners $(\eta_{n})_{n}$ if $(\boldsymbol{u}_{n},\nabla L_{n},P_{n},\eta_{n},\beta)$ are $(\delta_{n},\eta_{n})$-localizable with localizing sequence $(E_{K})_{K}$, and furthermore there exist locally Lipschitz functions $\tilde{\boldsymbol{h}}:\mathbb{R}^{k}\times\mathbb{R}\rightarrow\mathbb{R}^{k}$ and $\boldsymbol{\tilde{\Sigma}}:\mathbb{R}^{k}\to\mathbb{R}^{k\times k}$, such that
$\sup_{x\in\boldsymbol{u}_{n}^{-1}\left(E_{K}\right)}\left\|\left(-\frac{1}{1-\beta}\tilde{\mathcal{A}}_{n}+\frac{\delta_{n}}{(1-\beta)^{2}}\tilde{\mathcal{L}}_{n}\right)\boldsymbol{u}_{n}(x)-\tilde{\boldsymbol{h}}\left(\beta,\boldsymbol{u}_{n}(x)\right)\right\|\rightarrow 0$
$\sup_{x\in\boldsymbol{u}_{n}^{-1}(E_{K})}\|\delta_{n}J_{n}\tilde{V}J_{n}^{T}-\tilde{\boldsymbol{\Sigma}}(\boldsymbol{u}_{n}(x))\|\to 0$
Theorem 5.3.
Let $(X_{\ell}^{\delta_{n}})_{\ell}$ be SGD initialized from $X_{0}\sim\mu_{n}$ for $\mu_{n}\in\mathscr{M}_{1}\left(\mathbb{R}^{p_{n}}\right)$ with learning rate $\delta_{n}$, momentum rate $\beta\in[0,1)$ and pre-conditioner $\eta_{n}(x,y)$ for the loss $L_{n}(\cdot,\cdot)$ and data distribution $P_{n}$. For a family of summary statistics $\boldsymbol{u}_{n}=\left(u_{i}^{n}\right)_{i=1}^{k}$, let $\left(\boldsymbol{u}_{n}(t)\right)_{t}$ be the linear interpolation of $(\boldsymbol{u}_{n}(X_{\lfloor t\delta_{n}^{-1}\rfloor}^{\delta_{n}}))_{t}$.
Suppose that $\boldsymbol{u}_{n}$ are asymptotically closable with learning rate $\delta_{n}$, momentum rate $\beta$, preconditioners $(\eta_{n})_{n}$, effective drift $\tilde{\boldsymbol{h}}$, and effective volatility $\tilde{\boldsymbol{\Sigma}}$, and that the pushforward of the initial data has $\left(\boldsymbol{u}_{n}\right)_{*}\mu_{n}\rightarrow\nu$ weakly for some $\nu\in\mathscr{M}_{1}\left(\mathbb{R}^{k}\right)$. Then $\left(\boldsymbol{u}_{n}(t)\right)_{t}\rightarrow\left(\boldsymbol{u}_{t}\right)_{t}$ weakly as $n\rightarrow\infty$, where $\boldsymbol{u}_{t}$ solves:
$\text{d}\boldsymbol{u}_{t}=\tilde{\boldsymbol{h}}(\beta,\boldsymbol{u}_{t})\,\text{d}t+\frac{1}{1-\beta}\sqrt{\tilde{\boldsymbol{\Sigma}}\left(\boldsymbol{u}_{t}\right)}\,\text{d}\mathbf{B}_{t}$
This result is essentially identical to Theorem 2.3 under the revised constructions of $\nabla\Phi$, $\nabla H$, to $\nabla\tilde{\Phi}$ and $\nabla\tilde{H}$. That is, by augmenting the loss gradient with the preconditioner to obtain the coupled gradient $\eta_{n}\nabla L$, we can leverage the same mechanics as those for SGD-M to immediately obtain scaling limits for SGD-M equipped with Markovian scalar preconditioners. To understand why this holds, we note that $\nabla L_{n}$ is treated as an arbitrary vector field satisfying specific regularity conditions in the proof of Theorem 2.3. The fact that $\nabla L_{n}$ is the gradient is irrelevant, and in principle can be replaced with any arbitrary estimator satisfying Definitions 5.1, 5.2. In this case, we replace $\nabla L$ with the “estimator” $\eta\nabla L$.
@@
Let $\tau_{K}^{n}$ denote the exit time of the interpolated process $\boldsymbol{u}_{n}(t)$ from the compact set $E_{K}^{n}$. Furthermore, let $E^{*}_{K,n}:=\boldsymbol{u}_{n}^{-1}(E^{n}_{K})$ denote the corresponding pre-image, and $L^{\infty}_{K,n}:=L^{\infty}(E^{*}_{K,n})$ the associated space of localized bounded functions. The subsequent analysis relies on several technical lemmas, which hold under the aysmptotic-closability and - $\delta_{n}$-localizability assumptions detailed in Definition 2.1, 2.2. We assume $\boldsymbol{p}_{0}=0$ unless stated otherwise.
@@
By consolidating Lemmas 6.5, 6.6, and 6.7, we construct the auxiliary process $\boldsymbol{z}_{n}$ as follows. Consider the $j$-th component $z^{j}$, the discrete-time process is defined by the Doob decomposition:
@@
The remainder of the proof closely follows [5, Theorem 2.3]; we provide the details for completeness. Let $\boldsymbol{z}_{n}(s)$ be the linear interpolation of the discrete process $(\boldsymbol{z}_{\lfloor s/\delta\rfloor})$ given in (16). We decompose this process as
@@
The summation ranges over $\ell$ from $\lfloor s/\delta\rfloor+1$ to $\lfloor t/\delta\rfloor$, stopped at $\tau_{K}^{n}/\delta$. The bound follows from the continuity of $\boldsymbol{h}$ as stipulated by asymptotic closability, i.e. Definition 2.2.
@@
By asymptotic-closability, we ensure that as $n\to\infty$,
@@
By Itô’s formula for continuous semimartingales [51, Theorem 6.3.4], for any $f\in C_{0}^{\infty}(\mathbb{R}^{k})$, the process $f(\boldsymbol{z}^{K}_{t})-\int_{0}^{t}\mathsf{L}f(\boldsymbol{z}^{K}_{s})ds$ is a martingale, where $\mathsf{L}$ is the infinitesimal generator
@@
stopped at the exit time $\tau_{K}$. By employing a standard localization argument (see, e.g. [51, Chapter 11]), we conclude that every limit point $\boldsymbol{z}(t)$ of $\boldsymbol{z}_{n}(t)$ is a solution to the limiting SDE. ∎
6.1. Proof of Lemmas 6.1-6.3